OpenStudy (anonymous):

In triangle STU, SU = 18 and TU = 13. Find m

OpenStudy (anonymous):

is it a right triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

okay lets see hmmm

OpenStudy (anonymous):

this is literally an example of what I just asked like 10 minutes ago if you can find the missing angles in a right triangle if you have some of the langths

OpenStudy (anonymous):

By using the tan formula: tan (angle) = (opposite side) / (adjacent side) tan (T) = SU / TU T = tanh (SU / TU) Hope this helps.

OpenStudy (anonymous):

@dauspex how does tangent find angle measures though, I forget?

OpenStudy (anonymous):

UGH. I'm not getting any of this.

OpenStudy (anonymous):

tan (T) is actually a ratio of the sides of a triangle http://www.mathsisfun.com/sine-cosine-tangent.html @ehmuleeee suggest you take a read at the above site.

OpenStudy (anonymous):

Ok.